OpenStudy (anonymous):

g(x)=[g(x)=|cosx−1|. on[0,2π],what is the x value when g attains its maximum value?

OpenStudy (dumbcow):

range of cos is [-1,1] |-1-1| = |-2| = 2 |1-1| = 0 therefore range of g(x) is [0,2] max value is 2 which occurs when cos(x) = -1 From unit circle cos(x) = -1 when x = pi

OpenStudy (zarkon):

Jasleen...I guess you didn't believe me when I answered it before :)

OpenStudy (dumbcow):

oh haha, don't you hate that

OpenStudy (zarkon):

yep

OpenStudy (anonymous):

no i just wanted to see how its done!!

OpenStudy (anonymous):

and i have one more question that i need help with! f(x) =x^3+4x^2+4x. FIND the zeros of f.